OpenStudy (anonymous):

lim (1+1/x)^x x-> The answer is e but I do not know how to get there

OpenStudy (amistre64):

well, that is the definition for 'e' :)

OpenStudy (anonymous):

I do not understand what you mean

OpenStudy (anonymous):

sorry I did not mean to repeat

OpenStudy (amistre64):

i mean that 'e' is defined as the limit as x -> inf; (1+1/x)^x

OpenStudy (amistre64):

try using the definition of a derivative..

OpenStudy (amistre64):

f(x+h) - f(x) lim h->inf; ----------- h

OpenStudy (anonymous):

ok, I will try it

OpenStudy (anonymous):

one definition is it is the solution to \[\int^x_1 \frac{1}{t}\,dt = 1\]

OpenStudy (anonymous):

@amistre64: Its not right, the definition is \[h \rightarrow 0\] As to make the secant a tangent!

OpenStudy (amistre64):

lol...same diff :)

OpenStudy (anonymous):

Ok, thanks amistre64, but I am not grasping the point

OpenStudy (anonymous):

What do you mean ami

OpenStudy (amistre64):

i mean.... i might be wrong ;)

OpenStudy (amistre64):

what math is this for hix?

OpenStudy (anonymous):

calculus I

OpenStudy (anonymous):

another definition of e is \[e= \lim_{x\to\infty} (1+\frac{1}{x})^x\] in which case the answer is immediate.

OpenStudy (amistre64):

well, calc1 at this semester should be talking about limits right?

OpenStudy (anonymous):

i have a final tomorrow

OpenStudy (anonymous):

or you can take the log, compute the limit, get 1, and then 'exponentiate' to get e

OpenStudy (amistre64):

hmm.... my finals were 2 weeks ago; just starting the summer semester here

OpenStudy (anonymous):

I have to go to class, but please post anything that can help, and I will check it later. Thank you all very much. see ya

OpenStudy (anonymous):

\[lim_{x\to\infty} (1+\frac{1}{x})^x\] \[ln(1+\frac{1}{x})^x = x ln(1+\frac{1}{x})\] which is of the form \[\infty \times 0\] rewrite as \[\frac{ln(1+\frac{1}{x})}{\frac{1}{x}}\]

OpenStudy (anonymous):

Then use l'hopital. But this is a lot of unnecessary work since you have the definition of e to begin with.

OpenStudy (anonymous):

Let \[y = \lim_{x\to\infty}(1+1/x)^x\] we have to put it in indeterminate form in order to employ l'hopital's rule so we take natural log of both sides \[y = \lim_{x\to\infty}(1+1/x)^x \rightarrow \ln{y} = \lim_{x\to\infty}\ln((1+1/x)^x) \] Using log rule of exponents we have \[\lim_{x\to\infty}x*\ln((1+1/x))\rightarrow \infty-0\] Rewriting the whole expression \[\lim_{x\to\infty}\ln((1+1/x))/(1/x)\rightarrow 0/0 \] which is indeterminate, using L'Hopital's rule we have \[\lim_{x\to\infty}(-x/(x^2(x+1))/(-1/x^2)\rightarrow \lim_{x\to\infty} x/(x+1) \rightarrow \infty/\infty \] which permits us to use L'Hopital's rule again so we have \[\ln{y} = \lim_{x\to\infty}(1/1) = 1\] Raising both sides to e yields \[y = e^1=e\] Done.